# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

### v3

@@ 3 The Approach
Our proposed debiasing approach is schematically depicted in Figure 2. In this section, we provide at first a general description of the problem setting (Sec.  3.1), and then, we illustrate in detail DDB’s two main components, which include bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.3).
@@ 3.1 Problem Setting
Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution. Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e.possible biases such as context, appearance, acquisition noise, viewpoint, etc.. In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases. In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting [13] In this context, biased and unbiased samples equivalently refer to bias-aligned and bias conflicting samples.. If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization. This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 31]), which is dominating over the true target distribution semantics.
It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12]. For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in [12, 13]. While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
@@ 3.2 Diffusing the Bias
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [32]. Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\bm{\epsilon}_{\theta}$:
@@
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics. Classifier-Free Guidance (CFG) [33] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
@@ 3.3.1 Recipe I: 2-step debiasing
Figure 3: Overview of Recipe I’s 2-step debiasing approach.
The adopted 2-step approach consists in 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) apply a bias supervised method to train a debiased target model for classification. For the latter, we use the group DRO algorithm [8] (G-DRO) as a proven technique for the pure debiasing step. In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right)$ (3)
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
@@ 3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] auxiliary intentionally-biased models. Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method [6]. LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4. Such loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
Figure 4: Overview of Recipe II’s end-to-end debiasing approach.
@@ 4 Experiments
In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis.
In the Supplementary Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
@@ 4.1 Datasets
Our experiments utilize five distinct datasets: Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), Biased Action Recognition (BAR), and ImageNet-9/ImageNet-A. Waterbirds  [8] is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). It has 4,795 training images, 1,199 images for validation, and 5,794 for testing. BFFHQ [4] builds upon FFHQ [34] by introducing demographic biases for face recognition, while BAR  [6] is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts. For these two last datasets, we specifically refer to their original versions introduced respectively in [6] and [4]: BFFHQ has a total of 19,200 training images where  ( $y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$). Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes. BAR presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing. Finally, ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions. Originally introduced by Hendrycks et al. [35] to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 [36], a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases [36]. As such, both our CDPM and Bias Amplifier are trained on ImageNet-9. Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of [36, 1, 31]
4.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels. Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, and ImageNet-9 respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [37] scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 [38] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer [39]. For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips. For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds [6, 4, 23]. In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
@@ 4.3.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 5), to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample. In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [40]. We estimate a caption for each image per class and count the frequency of every word. Suppl. Material (Sec. A.E) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced. The same behavior also emerges in other datasets used in this study, where a consistent description of the bias attributes can be noted, as reported in Sec. A.E. These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
@@ 4.3.2 Classification accuracy
Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5]. For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20]. BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It’s worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
@@
[…]
CDvG+LfF [29] $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
@@
Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () indicate that the method is bias-unsupervised. We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1). Notably, in Waterbirds, DDB’s WGA is higher than G-DRO supervised (91.56% vs. 91.40% of G-DRO [8]) and all the other unsupervised methods’ ones. DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to ETF-Debias [43] (73.60%), the best state-of-the-art results to date, and provides a boost of 7.45% over the recent DeNetDM [45].

### v4

@@ 3 The Approach
Our proposed debiasing approach is schematically depicted in Figure 1. In this Section, we first describe the general problem setting (Sec. 3.1), then detail DDB’s two main components: bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.3).
@@ 3.1 Problem Setting
Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution. Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e., possible biases such as context, appearance, acquisition noise, viewpoint, etc.. In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not present the same biases. In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting kim2024training . If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization. This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity kim2021biaswap ; Sagawa*2020Distributionally ; nahon2023mining ), which is dominating over the true target distribution semantics.
It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation d2024openbias . For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in d2024openbias ; kim2024training . While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature nam2020learning ), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
@@ 3.2 Diffusing the Bias
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps DBLP:conf/nips/HoJA20 . Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\bm{\epsilon}_{\theta}$:
@@
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics. Classifier-Free Guidance (CFG) DBLP:journals/corr/abs-2207-12598 introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
@@ 3.3 DDB: Bias Amplifier and Model Debiasing
(a) Overview of Recipe I’s 2-step debiasing approach.
(b) Overview of Recipe II’s end-to-end debiasing approach.
Figure 2: Comparison of two debiasing strategies: Recipe I and Recipe II.
@@ 3.3.1 Recipe I: 2-step debiasing
The adopted 2-step approach consists of 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) applying a bias supervised method to train a debiased target model for classification. For the latter, we use the group DRO algorithm Sagawa*2020Distributionally (G-DRO) as a proven technique for the pure debiasing step. In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering sohoni2020no or anomaly detection pastore2024lookingmodeldebiasinglens , we adopt a simple heuristic based on the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right),$ (3)
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 2(a).
@@ 3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one nam2020learning or more NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 auxiliary intentionally-biased models. Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method nam2020learning . LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 2(b). Such a loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
@@ 4 Experiments
In this Section, we report the description of the benchmark datasets employed, followed by a discussion of the obtained results and a thorough ablation analysis.
In the Appendix, we report the general implementation (Sec. A.2) and training details, including the implementation details of Recipe I (Sec. A.3) and Recipe II (Sec. A.4), and more examples of the generated images (Sec. A.6).
@@ 4.1 Datasets
Our experiments utilize six distinct datasets: three typical benchmark datasets, including Waterbirds Sagawa*2020Distributionally , Biased Flickr-Faces-HQ (BFFHQ) kim2021biaswap , and Biased Action Recognition (BAR), taken from the original version as introduced in nam2020learning . On top of these benchmark datasets, we are interested in evaluating our approach in challenging scenarios involving real-world images with natural biases and datasets where multiple shortcuts are present. For this reason, we include ImageNet-9 bahng2019rebias /ImageNet-A hendrycks2021natural , and UrbanCars li2023whac in our experiments. More dataset details are in Sec. A.1.
@@ 4.2.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 3) to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming the conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by three independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample, where possible. Specifically, for Waterbirds, the reported average bias-aligned proportion of synthetic generated images is $98.35\%$, in BAR it is $99.20\%$, and $99.70\%$ is obtained for BFFHQ. All these values are higher than the known proportions of bias in the original datasets. In addition, focusing on Waterbirds, we also evaluated the generated image content by adopting an image captioner, specifically BLIP-2 DBLP:conf/icml/0008LSH23 . We estimate a caption for each image per class and count the frequency of every word. Supplementary Material (Sec. A.7) shows a histogram of the detected keywords, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
@@ 4.2.2 Classification accuracy
Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in Sagawa*2020Distributionally ; Lee_Park_Kim_Lee_Choi_Choo_2023 . For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following kim2021biaswap ; li2022discover . BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It is worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
@@
CDvG+LfF jung2023fighting $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
@@
Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1). Notably, in Waterbirds, DDB’s WGA is higher than all the other state-of-the-art unsupervised methods’ accuracies, as well as of that of the supervised G-DRO Sagawa*2020Distributionally (91.56% vs. 91.40% of G-DRO). DDB-I is providing slightly worse performance than DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN li2022discover ), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to the state-of-the-art ETF-Debias wang2024navigate (73.60%).
Method ID.ACC $\uparrow$ BG-Gap $\downarrow$ CoObj-Gap $\downarrow$ BG-CoObj-Gap $\downarrow$
GroupDRO Sagawa*2020Distributionally $91.60$ $10.90$ $3.60$ $16.40$
LLE li2023whac $96.70$ $2.10$ $2.70$ $5.90$
LfF nam2020learning $97.20$ $11.60$ $18.40$ $63.20$
JTT pmlr-v139-liu21f_JTT $95.90$ $8.10$ $13.30$ $40.10$
EIIL creager2021environment $95.50$ $4.20$ $24.70$ $44.90$
DebiAN li2022discover $\underline{98.00}$ $14.90$ $10.50$ $69.00$
[…]
Table 2: Comparison of deep learning methods on the UrbanCars biased dataset. ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better).
For UrbanCars (see Table 2), presenting multiple biases, we adopted the same metrics of li2023whac (see Suppl. Material, Sec. A.1). DDB-I shows overall the most balanced performance across bias-conflicting samples, reporting an improvement over LLE li2023whac of $0.25\%$ for BG-Gap, $2.18\%$ on CoObj-Gap, and $5.78\%$ for BG-CoObj-Gap. DDB-II provides a significant improvement over the original LfF, but struggles more with the background-conflicting samples than DDB-I.